UT 20230905: Restart Night#4



Night Info

For the night of: Mon, 2023 Sep 4 ---> Tue, 2023 Sep 5
Local midnight = 2023 Sep 5,  7 hr UT, or JD 2460192.792
Local Mean Sidereal Time at midnight =  22 36 58.7

Sunset ( 1000 m horizon):   18 45 MST; Sunrise:   5 52 MST
Evening twilight:  20 05 MST;  LMST at evening twilight:  18 41
Morning twilight:   4 32 MST;  LMST at morning twilight:   3 10
12-degr twilight: 19 35 MST -->  5 02 MST; night center:   0 19 MST

Moonrise:  21 38 MST   
Moon at civil midnight: illuminated fraction 0.663
1.6 days before last quarter, RA and dec:   3 22 19, +20 21.3

Weather:


Staff

  • OSA: Josh Williams
  • ISA/SO: Andrew Cardwell
  • ECD Coordinator: Jenny Power

Night Plan

On-Sky Plan

MODS Alignment
  1. MODS Sky flats
    • 3Kx3K imaging flats in all filters ugriz
  2. DONE AGw5/6 Alignment (MODS1/2)
  3. Field Aberration LDG, RDG: Requires Seeing <1.2"
  4. Rotator Centers on MODS detectors
    • Scripts and readme located in: /home/lbto/Service_Observing/Restart_2023/MODS/RotCen
    • ~30 minutes. 10 for the initial slew+collimation and then about 5 min per PA (~3 for the script, but add time for rotator slews)
    • Run the sequence:
      • active0 to collimate, then
      • track0
      • track180
      • track90
      • track270
  5. Rotator Zero Points
    • Scripts and MMS files are located: /home/lbto/Service_Observing/Restart_2023/MODS/RZP
    • These 3 masks cover the night. Try to align 2 masks. ~15 min each.
      • 20p25
      • DONE 00p25
      • 04p25
  6. DONE Transform
    • MODS AGW Transform data - <1.2" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky) - must do RotZP First
    • Low priority if RotZP unchanged
    • Dither up and down a minimum of 15 cycles with pointing mask
    • Field and scripts located in MODS Restart OT and scripts located with Mask adjusted to pointing mask in:
      /nfs/192.168.39.20/homes/lbto/Service_Observing/Restart_2023/MODS/Transform
    • If RZP has changed:
      • Update transform from measurements
        • Files: agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
        • stop AGW services (mod1 stop agw, mods2 stop agw)
        • new file copied to ~/Config/MODSN/ and renamed agw.ini
        • Restart the AGW services (mods1 start agw, mods2 start agw)
      • Collect transform data for verification
      • Measure Rot ZP with MODS MOS masks
      • Dither Test M52 or M71
      • Confirmation new transform is acceptable
  7. Photometric Standards
    • Available first half: Need to check nimgs in scripts (times based on scripts in Restart_2022).
      • M71 : twilight - 8UT: ~30 min (21 for the script and the extra for slew+collim)
        • dual (8 min)
        • red (7 min)
        • blue (6 min)
      • M92: twilight - 6 UT ~40 min
        • dual (12 min)
        • red (12 min)
        • blue (10 min)
      • M13: twilight - 5 UT
  8. Spectrophotometric Standards
    • BD+33 until 04:30UT, BD+28 until 10UT, Feige110 6-10UT, G191-B2B from 9:30UT to twilight. Try to get Feige110 and/or G191-B2B which means standards are 2nd half of night.
    • If Clear, 2-3 spectra/mode
      For 3 spectra/mode, this will take 1hr/star
    • If not clear, 1 spectrum/mode
      One spectrum/mode will take 36 min
  9. Science-like script
    • Similar to transform data collection so optional
    • Science-Like with Dual Grating: NGC 40 - 00hr +72 (generate from Restart OT)
    • Dual Imaging NGC6764_ugriz.xml (MODS_2023/MODS_2023_workspace) 19hr +50d (seeing <1.2")
      • May need to adjust exposure times

LBC Checkout
  1. Twilight flats: Try to obtain flats through all 6 B and 8 R filters during the course of restart
  2. Focal Plane Tilt Check: focal plane tilt with superfoc sequences at PA=0, PA=180, PA=0 /lbt/lbto/lbc/Calib_OBs/FOCUS/(OBNAME)
  3. TMS testing: Run TMS with the following This will check:
    1. Standard Fields: If photometric, observe at least one photometric standard star field from each set (also tests dithering, but not guiding):
      1. Clem & Landolt
        • LBCB: UBV
        • LBCR: VRI
      2. sdssY
        • LBCB:gr (+Us as filler)
        • LBCR: rizY
    2. Observe with guiding and dithering: There are a few science-like OBs in the checkout OT
      1. BlueSnowball (23h42d)
      2. LittleSnoball (1hr51deg)

Closed Dome Plan

  1. SX Flat optimization with LBTI WFS
    1. Configure the telescope for LBTI (can not be done in parallel to moving telescope or swingarm tasks)
    2. Requires stow mins in, on oil, dark chamber, argos source deployed
    3. Acquire the argos source with the LBTI WFS, close the loop. Save a new flat. Rinse and repeat
  2. MODS reference Calibrations: Contingent on repairs to MODS1 quadcell and completion of tasks 1 & 2
    • Telescope at zenith, lights off for all but the lamps. Take these as able, but try to get through the (1) BIases, (2) Comparison lamps (grating), (3) Slitless pixel flats and (4) Imaging flats and then the dual grating slit flats.
    1. Slit flats:
      • Dual
        • 2.4" - 18 min
        • 5" - 18 min
        • 0.3" -
      • Red (67, w/o 0.3")
        • 0.3"
      • Blue (81, w/o 0.3")
        • 0.3" - 15 min
    2. Comparison lamps (prism = prlamps)
      • 23 min total - good check on whether all the lamps are working.
        • dual - 8 min
        • red - 7 min
        • blue - 7 min


Night log

https://docs.google.com/document/d/1ba-P4n2tC7b50dciYWnaOuHPYrPJ0Ty8IEYcXu8644M/edit?usp=sharing

-- JenniferPower - 04 Sep 2023
Topic revision: r6 - 05 Sep 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback